OpenStudy (anonymous):

MEDAL + Fan Word problems During the​ women's basketball championship​ game, team A scored 4 more points than team B.​ Together, both teams scored a total of 152 points. How many points did the Champion team A score during this​ game?

OpenStudy (anonymous):

I believe the answer team A scored 78 points, because 152-4=148 148/2=74 74+4=78

OpenStudy (anonymous):

why did you subtract first ?

OpenStudy (anonymous):

Because team A scored 4 more points than team B. So you have to subtract those for extra points.

OpenStudy (anonymous):

four*

OpenStudy (anonymous):

Alright I think I understand.

OpenStudy (denonakavro):

Lets say team B is x and team A is x+4 and total is equal to 152. Our equation would be x+x+4=152 x+x=152-4 2x=148 x=74 Team B x+4= 78 Team A

OpenStudy (anonymous):

A video game system and several games are sold for $548. The cost of the games is 3 times as much as the cost of the system. Find the cost of the system and the cost of the games. So, this one would be the separation between the System and the games. The games would be 3x-x= 548?

OpenStudy (anonymous):

@denonakavro + @ayeohannah

OpenStudy (anonymous):

I could totally be wrong, but I believe the cost of the games is 411 and the system is 137. I say that because 548/4= 137 and if you multiply 3 by 137 you have the total of the games (411) and that leaves 137 to be the total of the system... I could very well be wrong but I think that's how this one is solved.

OpenStudy (anonymous):

I have no idea. I can never do word problems. So id rather go with that than mine. so thank you.